OpenStudy (anonymous):

Solve this equation over the interval [0,2π) 1. 2cosx+4=5

OpenStudy (tkhunny):

Can you solve this for z? 2z + 4 = 5

OpenStudy (mtalhahassan2):

2z=5-4 z= 1/2